OpenStudy (anonymous):

I need help wit dis one.... -15,-9,-3,3_,_,_

geerky42 (geerky42):

-15 + 6 = -9 -9 + 6 = -3 -3 + 6 = 3 3 + 6 = ???

OpenStudy (eli_moses):

-15, -9, -3, 3, 9, 15

OpenStudy (eli_moses):

3, 9, 15, 21

OpenStudy (dumbcow):

"wit dis one" ?? i think you meant to say "with this one" :)